Find an equation of the line. Write the equation using function notation.

Through (3,−6)​; perpendicular to 4y=x


1
Expert's answer
2022-02-15T17:04:19-0500

"y= \frac{1}{4}x"

Gradient of the line "=-(\frac{4}{1})=-4"


"y=-4x+c"

When "x=3,\>y=-6"

"-6=-4(3)+c"

"c=6"

"f(x)=-4x+6"